"\intop"24 ex cosxdx=?


1
Expert's answer
2022-01-12T18:29:16-0500
"\int e^x\cos xdx=e^x\sin x-\int e^x\sin xdx"

"=e^x\sin x-(-e^x\cos x+\int e^x\cos xdx)"

"=e^x\sin x+e^x\cos x-\int e^x\cos x dx"

"\int e^x\cos xdx=\dfrac{1}{2}(e^x\sin x+e^x\cos x)+C"

"\displaystyle\int_{2}^{4}e^x\cos xdx=[\dfrac{1}{2}(e^x\sin x+e^x\cos x)]\begin{matrix}
 4 \\
 2
\end{matrix}"

"=\dfrac{1}{2}(e^4\sin 4+e^4\cos 4-e^2\sin 2-e^2\cos 2)"